# SLED: A Speculative LLM Decoding Framework for Efficient Edge Serving

### v1

@@ Abstract
Regardless the advancements in device capabilities, efficient inferencing advanced large language models (LLMs) at the edge remains challenging due to limited device memory and power constraints. Existing strategies, such as aggressive quantization, pruning, or remote inference, trade accuracy for efficiency or lead to substantial cost burdens. This position paper introduces a new approach that leverages speculative decoding, previously viewed primarily as a decoding acceleration technique for autoregressive generation of LLMs, as a promising approach specifically adapted for edge computing by orchestrating computation across heterogeneous devices. We propose SLED, a method that allows lightweight edge devices to draft multiple candidate tokens locally using diverse draft models, while a single, shared edge server efficiently batches and verifies the tokens utilizing a more precise target model. This approach supports device heterogeneity and reduces server-side memory footprint by avoiding the need to deploy multiple target models. Our initial experiments with Jetson Orin Nano, Raspberry Pi 5, and an RTX 6000 edge server indicate substantial benefits: significantly reduced latency, improved energy efficiency, and increased concurrent inference sessions, all without sacrificing model accuracy.
@@ I Introduction
Deploying LLMs at the edge, closer to data sources and end-users, offers significant advantages including reduced latency, enhanced privacy, and lower bandwidth consumption[6]. Nevertheless, edge environments, characterized by limited memory, processing power, and energy budgets, present formidable obstacles to efficient LLM inference. Existing strategies to address these limitations include aggressive model compression techniques such as quantization (reducing numerical precision) [7, 8], pruning (removing redundant parameters) [9, 10], and knowledge distillation (training smaller models to mimic larger ones) [11]. Other approaches involve distributed inference, where model layers are partitioned across multiple devices or between edge and cloud, or full remote inference, where the entire computation is offloaded to a powerful central server [12, 13, 14, 15]. While these methods offer some relief, they often come with trade-offs: compression can sacrifice model accuracy, distributed inference introduces synchronization overheads, and remote inference negates the benefits of edge deployment, incurring latency and communication costs.
This position paper introduces SLED, a novel approach that re-imagines speculative decoding as a paradigm specifically tailored for efficient LLM inference at the edge, by intelligently orchestrating computation across heterogeneous devices. SLED enables lightweight edge devices (e.g., IoT sensors, mobile phones, single-board computers) to locally draft multiple candidate tokens using diverse, small, and efficient draft models. Concurrently, a single, shared edge server, equipped with a more precise target model, efficiently batches and verifies these drafted tokens. This disaggregated architecture uniquely supports device heterogeneity, as edge devices can utilize a variety of draft models best suited to their capabilities, and crucially, it drastically reduces the server-side memory footprint by eliminating the need to deploy multiple, potentially large, target models. Our core motivation stems from the insight that significant GPU resources are often ”wasted” on generating ”easy” tokens in autoregressive decoding; SLED instead leverages the server’s power primarily for the computationally intensive task of verifying tokens, thus maximizing resource utilization.
@@
  • We introduce a disaggregated architecture that partitions the speculative decoding process across lightweight edge devices and a powerful, shared edge server, optimizing resource utilization and minimizing server-side memory footprint.
  • We demonstrate through empirical evaluation the substantial benefits of SLED in terms of reduced latency, improved energy efficiency, and higher generation quality on diverse edge hardware.
The remainder of this paper is organized as follows: Section II reviews existing work in LLM inference, edge computing, and speculative decoding. Section III details the architectural design and key components of SLED. Section IV presents our experimental setup and discusses the evaluation results. Finally, Section V and VI concludes the paper and outlines future research directions.
@@ II Related Work
Model Compression and Lightweight Architectures. To enable LLMs to run on resource-constrained devices, significant efforts have been directed towards model compression. Quantization reduces the numerical precision of model parameters and activations (e.g., from 32-bit floating point to 8-bit integers or even binary) to decrease memory footprint and accelerate computation [16, 17, 18]. While highly effective, aggressive quantization can lead to a noticeable drop in model accuracy. Pruning identifies and removes redundant connections or neurons from the neural network without significant performance loss, resulting in sparser and smaller models [19]. Knowledge distillation involves training a smaller ”student” model to mimic the behavior of a larger, more powerful ”teacher” model, thereby transferring knowledge and achieving comparable performance with a significantly smaller footprint [11]. Beyond these optimization techniques, research has also focused on designing inherently lightweight transformer architectures that are more efficient from the ground up, such as MobileBERT [20], Mamba [21] or other compact variants, often by optimizing attention mechanisms or reducing the number of layers and hidden dimensions. Despite their advantages in reducing model size and computational demands, a common limitation of these model compression techniques is the inherent trade-off with model quality; aggressive compression often leads to a measurable decrease in accuracy or generation fidelity compared to their full-sized counterparts.
@@
Edge-Cloud/Server Offloading and Distributed Inference. Another line of research focuses on distributing LLM computation across multiple devices or partitioning tasks between edge and cloud/server infrastructure. Model partitioning schemes divide a large LLM into smaller sub-models, with different parts executed on different devices (e.g., some layers on the edge, others on a server) [12, 22, 15]. This often involves pipeline parallelism or tensor parallelism techniques, where different stages or segments of the model’s computation are assigned to different devices. While this allows larger models to run on resource-constrained setups, it introduces communication overheads and synchronization challenges, particularly for heterogeneous hardware and varying network conditions. Edge-cloud offloading dynamically decides which parts of the inference task should be performed locally at the edge and which should be offloaded to a more powerful cloud server, often based on real-time resource availability, network bandwidth, and latency requirements [12]. These methods aim to balance the benefits of edge processing with the computational power of the cloud, but often require sophisticated orchestration and robust connectivity.
Pure Remote Inference. The most straightforward approach for deploying large LLMs is pure remote inference, where the entire model resides and executes on a powerful central server or in the cloud. User queries are sent to this central server, which performs the inference and returns the results. This approach leverages the centralized computational resources, ensuring that the most powerful and accurate LLMs can be utilized without local device constraints. It simplifies deployment and updates, as the model only needs to be managed in one location. However, pure remote inference incurs significant network latency, especially for interactive applications, and can lead to high bandwidth costs. Crucially, the standard decoding process in remote inference is often autoregressive, generating one token at a time, which can be memory-intensive due to large key-value caches and lead to resource under-utilization on powerful servers. The cost associated with cloud GPU instances for continuous, often under-utilized, inference also presents a substantial economic burden, especially for high-throughput scenarios.
@@
SLED directly addresses these limitations by fundamentally altering the inference paradigm. Instead of autoregressive token generation on the powerful central server, SLED offloads the preliminary token drafting to lightweight edge devices. This allows the central server to focus its considerable resources primarily on the more efficient and batchable task of verifying multiple drafted tokens. By doing so, SLED mitigates issues of sophisticated orchestration and significantly improves the utilization of expensive server-side GPU resources without sacrificing model accuracy, leading to a more cost-effective and scalable distributed LLM inference system.
@@ III SLED Design
The architecture of the proposed SLED framework is depicted in Fig.1. Within the defined service area, an edge infrastructure is composed of an edge server and $N$ heterogeneous edge devices, each equipped with their own lightweight large language models (LLMs) scaled according to individual computational and memory resource capacities. These edge devices are responsible for serving diverse LLM-based applications such as intelligent personal assistants, text generation, and semantic analysis, among other tasks.
In close proximity to the edge devices, typically located at facilities such as base stations, the edge server provides substantial computational capabilities, leveraging specialized hardware like Graphics Processing Units (GPUs) or Neural Processing Units (NPUs). On this edge server, a single, comprehensive target model is deployed, optimized for efficiently verifying the accuracy of draft tokens generated by the distributed edge devices.
Operationally, user-generated prompts, encompassing a wide array of task-specific requests, are initially received and tokenized locally by each edge device. Subsequently, the tokenized prompts, denoted as input sequences $p_{i}^{n}$ where $n\in\{1,2,\cdots,N\}$, are processed by local draft models to generate speculative tokens. These draft tokens are then transmitted to the edge server for validation. Upon completion of the verification step, the edge server communicates the results back to the respective edge devices, specifically identifying rejected token positions along with any necessary corrective tokens.
This drafting-verification workflow iteratively progresses, alternating between local speculation at the edge devices and centralized validation at the edge server, until the generated output reaches the predetermined desired length or the end-of-response token is encountered. This collaborative mechanism not only optimizes resource utilization by distributing computational tasks according to device capabilities but also significantly reduces latency and enhances overall efficiency in edge-based LLM serving scenarios.
@@ III-A Dynamic Drafting on Edge Devices
On edge devices, each verification cycle initiated by the edge server is preceded by the generation of multiple draft tokens. The acceptance rate of these draft tokens serves as a crucial indicator of their quality, directly influenced by the capabilities of the draft models utilized. A higher acceptance rate is desirable as it signifies fewer verification iterations and consequently reduces the computational burden on the costly target model, thereby mitigating communication overhead inherent in edge computing scenarios.
Previous studies [some_paper], corroborated by our preliminary experimental results, have established a correlation between the acceptance rate of draft tokens and their associated confidence scores derived from the output logits. As illustrated in Fig.2, draft tokens with higher confidence scores exhibit a significantly increased likelihood of acceptance by the target model. Further analysis also highlights additional statistical measures—such as entropy and margin of the output logits—as predictors for token acceptance probability, depicted in Fig.3.
Building upon these insights, we propose and implement a dynamic drafting mechanism on edge devices. This adaptive strategy modulates the speculative decoding length based on the real-time evaluation of token confidence scores. Formally, we introduce a threshold parameter, $c_{th}$, derived empirically, and define the decision-making process for triggering server verification for the $i$-th draft token $t_{s}^{i}$ in speculation round $s$ as follows:
$\begin{cases}c_{s}^{i}<c_{th},request\;verification\\ c_{s}^{i}\geq c_{th},generate\;another\;token\\ \end{cases}$ (1)
@@
Considering the inherent unreliability and fluctuating nature of network conditions in edge computing environments—such as variable round-trip time (RTT) and intermittent connectivity—we further enhance our system with an asynchronous decoding mechanism accompanied by a timeout protocol. Specifically, edge devices continue generating additional draft tokens using local lightweight language models (LLMs) concurrently while awaiting verification responses from the edge server. If a verification response confirms acceptance of all previously sent draft tokens, these locally generated tokens seamlessly transition into the draft token queue for subsequent verification cycles, thus significantly reducing idle wait times.
@@
Figure 3: The draft token acceptance rate with different logits entropy and margin is shown in Fig. 3. (a) Entropy vs. Acceptance Rate. (b) Margin vs. Acceptance Rate success ratio.
@@ IV Evaluation
In this section, we systematically evaluate the performance and efficiency of the proposed SLED framework through extensive simulations and measurements. We assess SLED’s efficacy compared to a centralized large language model (LLM) serving system across various metrics including throughput, computational resource utilization, energy efficiency, latency, and robustness under network uncertainties.
@@ IV-A Whole-System Token Generation Rate (WSTGR)
We first evaluate the Whole-System Token Generation Rate (WSTGR), which quantifies the productive output of the entire SLED system. Fig. 4 illustrates the WSTGR against the number of edge devices. We observe that, unlike the baseline centralized system, which saturates at approximately 40 devices due to computational bottlenecks, SLED scales nearly linearly up to 250 devices, significantly increasing overall system throughput. This scalability demonstrates that SLED effectively leverages distributed edge resources to enhance token generation capacity.
IV-B GPU-Server FLOP Load vs. Tokens Produced
Figure 5: Server-side GPU computational load relative to tokens produced, showing enhanced computational efficiency in the SLED system compared to the centralized baseline.
To further analyze computational efficiency, we compare server-side GPU utilization with the system-wide generated tokens. As depicted in Fig. 5, the centralized LLM serving system follows a linear trajectory, indicating each token requires a full forward pass on the GPU. Conversely, the SLED system demonstrates a sub-linear relationship, attributed to the selective verification process, thus reducing redundant computations. This indicates a substantial improvement in GPU resource utilization efficiency in the SLED framework.
IV-C Energy Consumption per Token
Figure 6: Comparative analysis of energy consumption per token across diverse hardware platforms, demonstrating the energy-efficiency advantages of the SLED framework.
We next assess energy efficiency across various hardware platforms by measuring energy per generated token. Table 6 lists these results across multiple hardware tiers and models. SLED consistently demonstrates superior energy efficiency compared to the remote inference baseline. Specifically, lower-power devices such as Jetson and Raspberry Pi exhibit significantly reduced energy consumption per token when employing local drafting and verification processes enabled by SLED, reflecting the advantage of computational offloading and reduced dependence on high-power edge servers.
IV-D Energy-Latency Pareto Front
@@
To illustrate the optimal trade-offs achieved by SLED, we evaluate the Pareto front of energy per token against the 95th percentile (P95) latency, as shown in Fig. 7. Compared to fully local inference on small models, SLED offers superior Pareto-optimal points by efficiently balancing latency and energy consumption. This demonstrates SLED’s capability to provide low-latency responses while minimizing energy overhead, crucial for latency-sensitive edge computing scenarios.
IV-E System Throughput vs. Server GPU Budget
Figure 8: Relationship between GPU hardware budget and system throughput, identifying cost-effective GPUs for optimized edge deployments.
Fig. 8 examines system throughput (WSTGR) in relation to different GPU hardware options based on their associated costs (normalized to the A100 GPU). Interestingly, while high-end GPUs like the A100 deliver the highest throughput, mid-range GPUs such as RTX 6000 Ada provide optimal cost-efficiency, producing competitive throughput at significantly lower total cost of ownership (TCO). This insight is critical for practical deployment scenarios, suggesting mid-tier GPUs might offer the best balance between cost and performance for edge-based LLM serving.
IV-F Robustness Under Network Uncertainty
Figure 9: Impact of RTT delay on communication rounds, illustrating SLED’s robustness through its adaptive rollback mechanism.
Lastly, we analyze the robustness of SLED against network uncertainties by measuring communication rounds required under varying round-trip time (RTT) delays (Fig. 9). SLED’s rollback mechanism effectively mitigates the impact of increased RTT delays by substituting verified results with locally generated tokens when network conditions degrade. This results in fewer communication rounds compared to the centralized baseline, highlighting SLED’s enhanced resilience and robustness in dynamic network environments.
Overall, the experimental evaluations underscore the significant advantages of SLED in distributed inference scenarios, including scalability, computational efficiency, energy savings, optimal latency-energy trade-offs, cost-effective resource utilization, and enhanced robustness against network fluctuations.

### v2

@@ Abstract
Regardless of the advancements in device capabilities, efficient inferencing advanced large language models (LLMs) at the edge remains challenging due to limited device memory and power constraints. Existing strategies, such as aggressive quantization, pruning, or remote inference, trade accuracy for efficiency or lead to substantial cost burdens. This position paper introduces a new approach that leverages speculative decoding, previously viewed primarily as a decoding acceleration technique for autoregressive generation of LLMs, as a promising approach specifically adapted for edge computing by orchestrating computation across heterogeneous devices. We propose SLED, a method that allows lightweight edge devices to draft multiple candidate tokens locally using diverse draft models, while a single, shared edge server efficiently batches and verifies the tokens utilizing a more precise target model. This approach supports device heterogeneity and reduces server-side memory footprint by avoiding the need to deploy multiple target models. Our initial experiments with Jetson Orin Nano, Raspberry Pi 4B/5, and an edge server equipped with 4 Nvidia A100 GPUs indicate substantial benefits: significantly increased system throughput, capacity, and better cost efficiency, all without sacrificing model accuracy.
@@ I Introduction
Deploying LLMs at the edge, closer to data sources and end-users, offers significant advantages including reduced latency, enhanced privacy, and lower bandwidth consumption[6]. Nevertheless, edge environments, characterized by limited memory, processing power, and energy budgets, present formidable obstacles to efficient LLM inference. Existing strategies to address these limitations include aggressive model compression techniques such as quantization  [7, 8], pruning [9, 10], and knowledge distillation  [11]. Other approaches involve distributed inference, where model layers are partitioned across multiple devices or between edge and cloud[12, 13], or full remote inference, where the entire computation is offloaded to a powerful central server [14, 15]. While these methods offer some relief, they often come with trade-offs: compression can sacrifice model accuracy, distributed inference introduces synchronization overheads and is incompatible with heterogeneous edge devices, and remote inference negates the benefits of edge deployment, incurring non-negligible costs.
This position paper introduces SLED, shown in Fig. 1, a novel approach that re-imagines speculative decoding as a paradigm specifically tailored for efficient LLM inference at the edge, by intelligently orchestrating computation across heterogeneous devices, as shown in Fig.2. Within the defined service area consisting of an edge server and $N$ heterogeneous edge devices, and each dege device is equipped with their own lightweight large language models (LLMs) scaled according to individual computational and memory resource capacities. These edge devices are responsible for serving diverse LLM-based applications such as intelligent personal assistants, text generation, and semantic analysis, among other tasks. Concurrently, a single, shared edge server, equipped with a more precise target model, efficiently batches and verifies these drafted tokens. This disaggregated architecture uniquely supports device heterogeneity, as edge devices can utilize a variety of draft models best suited to their capabilities, and crucially, it drastically reduces the server-side memory footprint by eliminating the need to deploy multiple, potentially large, target models to handle outputs from various draft models. Our core motivation stems from the insight that significant GPU resources are often ”wasted” on generating ”easy” tokens in autoregressive decoding; SLED instead leverages the server’s power primarily for the computationally intensive task of verifying tokens, thus maximizing system capacity, which is the number of supported devices. We compare the system capacity of SLED and centralized serving system with the same response rate but different device types. From Tab. I, we observe that compared with centralized LLM serving system for the edge, the proposed SLED is capable of increasing the system capacity by 2.6 to 2.9 times.
TABLE I: Capacity of SLED and centralized serving system
System Raspberry Pi 4b Raspberry Pi 5 Nvidia Jetson
SLED 18.30 5.24 19.53
Centralized serving 7.05 1.83 7.06
@@
  • In SLED, we propose and deploy the dynamic drafting scheme on edge devices. By dynamically requesting for verification according to the confidence score of draft model, the edge devices are able to avoid unnecessary verification, hence reducing the communication rounds and improve the utilization on the server.
  • We demonstrate through empirical evaluation the substantial benefits of SLED in terms of higher throughput, capacity and better cost-efficiency on diverse edge hardware.
The remainder of this paper is organized as follows: Section II reviews existing work in LLM inference for edge computing. Section III details the architectural design and key components of SLED. Section IV presents our experimental setup and discusses the evaluation results. Finally, Section V and VI concludes the paper and outlines future research directions.
@@ II Related Work
TABLE II: Comparison of related work; Edge-Serving: Does the system support edge computing?; Heterogenity: Is the heterogenity of edge devices considered in the system design?; Lossless: Whether does the system deliver LLM service without any performance degradation? Scalable Model: Is the system capable of scaling model according to conditions without too much overhead?
System Edge-Serving Heterogenity Lossless Scalable Model
EdgeShard[12]
Galaxy[13]
Orca[16]
vLLM[17]
FastServe[18]
AWQ[19]
MobileBERT[20]
SLED
@@
Model Compression and Lightweight Architectures. To enable LLMs to run on resource-constrained devices, significant efforts have been directed towards model compression. Quantization reduces the numerical precision of model parameters and activations to decrease memory footprint and accelerate computation [21, 22, 19]. Pruning identifies and removes redundant connections or neurons from the neural network without significant performance loss, resulting in sparser and smaller models [23]. Knowledge distillation involves training a smaller ”student” model to mimic the behavior of a larger ”teacher” model, thereby transferring knowledge and achieving comparable performance with a significantly smaller footprint [11]. Beyond these optimization techniques, research has also focused on designing inherently lightweight transformer architectures that are more efficient from the ground up, such as MobileBERT [20], Mamba [24] or other compact variants, often by optimizing attention mechanisms or reducing the number of layers and hidden dimensions. Despite their advantages in reducing model size and computational demands, a common limitation of these model compression techniques is the inherent trade-off with model quality; aggressive compression often leads to a measurable decrease in accuracy or generation fidelity compared to their full-sized counterparts.
@@
Edge-Cloud/Server Offloading and Distributed Inference. Another line of research focuses on distributing LLM computation across multiple devices or partitioning tasks between edge and cloud/server infrastructure. Model partitioning schemes divide a large LLM into smaller sub-models, with different parts executed on different devices [12, 25, 13]. For example, EdgeShard[12] partitioned LLM into shards and deploy on distributed devices to benefit from the collaboration among edge devices and cloud server. This often involves pipeline parallelism or tensor parallelism techniques, where different stages or segments of the model’s computation are assigned to different devices. While this allows larger models to run on resource-constrained setups, it introduces communication overheads and synchronization challenges, particularly for heterogeneous hardware and varying network conditions. Edge-cloud offloading dynamically decides which parts of the inference task should be performed locally at the edge and which should be offloaded to a more powerful cloud server, often based on real-time resource availability, network bandwidth, and latency requirements [12]. These methods aim to balance the benefits of edge processing with the computational power of the cloud, but often require sophisticated orchestration and robust connectivity.
Pure Remote Inference. Pure remote inference, where the entire large language model (LLM) resides on centralized data-center GPUs, represents a prevalent deployment paradigm due to its simplicity and centralized resource utilization. Recent research primarily focuses on resource efficiency and latency optimization. Kwon et al.[17] proposed vLLM with a PagedAttention allocator, significantly reducing KV-cache overhead and fragmentation, achieving up to 4× throughput improvement. Wu et al. [18] introduced FastServe, leveraging multi-level feedback queue scheduling and proactive KV-cache management to reduce tail latency by up to 31× at the 99th percentile. Rajbhandari et al. [26] developed DeepSpeed Inference, combining multiple parallelism strategies with NVMe and CPU off-loading, allowing inference of substantially larger models and reducing latency by up to 7.3×. Crucially, the standard decoding process in remote inference is often autoregressive, generating one token at a time, which can be memory-intensive due to large key-value caches and lead to resource under-utilization on powerful servers. Moreover, the cost associated with cloud GPU instances for continuous, often under-utilized, inference also presents a substantial economic burden, especially for high-throughput scenarios.
@@
Tab. II compares the SLED and related works that deliver LLM inference service or propose model variants for edge devices. As illustrated in Tab. II, SLED stands out as the only approach that enables lossless LLM inference for heterogeneous edge devices, while maintaining a collaborative design that can flexibly accommodate increasingly large models without requiring full replication at each endpoint. SLED directly addresses these limitations by fundamentally altering the inference paradigm. Instead of autoregressive token generation on the powerful central server, SLED offloads the preliminary token drafting to lightweight edge devices. This allows the central server to focus its considerable resources primarily on the more efficient and batchable task of verifying multiple drafted tokens. By doing so, SLED significantly improves the utilization of expensive server-side GPU resources without sacrificing model accuracy, leading to a more cost-effective and scalable distributed LLM inference system.
@@ III SLED Design
Fig. 2 shows the detailed structure and data flow of the SLED. In close proximity to the edge devices, typically located at facilities such as base stations, the edge server provides substantial computational capabilities, leveraging specialized hardware like Graphics Processing Units (GPUs) or Neural Processing Units (NPUs). On this edge server, a single, comprehensive target model is deployed, optimized for efficiently verifying the draft tokens generated by the distributed edge devices.
Operationally, user-generated prompts, encompassing a wide array of task-specific requests, are initially received and tokenized locally by each edge device. Subsequently, the tokenized prompts, denoted as input sequences $p_{i}^{n}$ where $n\in\{1,2,\ldots,N\}$, are processed by local draft models to generate speculative tokens. These draft tokens are then transmitted to the edge server for validation. Upon completion of the verification step, the edge server communicates the results back to the respective edge devices, specifically identifying rejected token positions along with any necessary corrective tokens.
This drafting-verification workflow iteratively progresses, alternating between local speculation at the edge devices and centralized validation at the edge server, until the generated output reaches the predetermined desired length or the end-of-response token is encountered. This collaborative mechanism not only optimizes resource utilization by distributing computational tasks according to device capabilities but also significantly reduces latency and enhances overall efficiency by transmitting tokens rather huge activations.
@@ III-A Dynamic Drafting on Edge Devices
On edge devices, each verification cycle is preceded by the generation of multiple draft tokens. The acceptance rate of these draft tokens serves as a crucial indicator of their quality, directly influenced by the capabilities of the draft models utilized. A higher acceptance rate is desirable as it signifies fewer verification iterations and consequently reduces the computational burden on the costly target model, thereby mitigating communication overhead inherent in edge computing scenarios.
Previous studies [27], corroborated by our preliminary experimental results, have established a correlation between the acceptance rate of draft tokens and their associated confidence scores derived from the output logits. As illustrated in Fig.3, draft tokens with higher confidence scores exhibit a significantly increased likelihood of acceptance by the target model.
Building upon this insight, we propose and implement a dynamic drafting mechanism on edge devices. This adaptive strategy modulates the speculative decoding length based on the real-time evaluation of token confidence scores. Formally, we introduce a threshold parameter, $c_{th}$, derived empirically, and define the decision-making process for triggering server verification for the $i$-th draft token $t_{s}^{i}$ in speculation round $s$ as follows:
$c_{s}^{i}\begin{cases}<c_{th},request\;verification\\ \geq c_{th},generate\;another\;token\\ \end{cases}$ (1)
@@
Considering the inherent unreliability and fluctuating nature of network conditions in edge computing environments, such as variable round-trip time (RTT) and intermittent connectivity, we further enhance our system with an asynchronous decoding mechanism accompanied by a timeout protocol. Specifically, edge devices continue generating additional draft tokens using local lightweight language models (LLMs) concurrently while awaiting verification responses from the edge server. If a verification response confirms acceptance of all previously sent draft tokens, these locally generated tokens seamlessly transition into the draft token queue for subsequent verification cycles, thus significantly reducing idle wait times.
@@ IV Evaluation
In this section, we systematically evaluate the performance and efficiency of the proposed SLED framework through extensive simulations and measurements. We assess SLED’s efficacy compared to a centralized LLM serving system which serves the decoding requests from edge device directly, and the edge-only inference system which generates all tokens locally on the devices across various metrics including throughput, system capacity, cost efficiency, and impact of speculative length on system capacity and the throughput.
To accurately simulate verification request workloads from edge devices utilizing speculative decoding, we adopt a Poisson-based modeling approach. Each edge device is considered an independent source of verification requests, with inter-arrival times following an exponential distribution. This modeling choice effectively captures the asynchronous and inherently stochastic nature of real-world device interactions, ensuring that the simulated workload closely mirrors realistic operational conditions. The device-specific request rate is derived directly from realistic device speculative decoding throughput, ensuring that the simulation’s temporal patterns closely align with practical speculative decoding workloads.
As for the device setting, we tested Raspberry pi 4b, 5 and Ndivia Jetson Orin Nano on the SLED system supported by 4 A100 GPUs.
@@ IV-A Whole-System Token Generation Rate (WSTGR)
We first evaluate the Whole-System Token Generation Rate (WSTGR), which is defined as the total number of tokens generated and verified by the entire inference system per second, and serves as a metric for the system’s overall productive output. Figure 4 presents the WSTGR as a function of the number of edge devices within the edge server’s service area. The verification workload model is derived from a Raspberry Pi 5 device running a LLaMA 3B model. Additionally, we evaluate two different target models (11B and 70B) on both the SLED system and a centralized inference system. As shown in Fig. 4, for both the 11B and 70B models, the WSTGR increases rapidly in the initial stages as batch size grows, due to the amortization of fixed GPU launch and driver overhead, and improved utilization of GPU cores. The proposed SLED system achieves higher overall throughput than the centralized serving system under identical conditions, including the same number of devices and target model. This observed scalability demonstrates that SLED effectively utilizes distributed edge resources to enhance the system’s token generation capacity.
The performance gain over the centralized serving system stems from the efficient and balanced distribution of computational tasks in the SLED system. Specifically, in SLED, simple token generation tasks can be handled by relatively small models [28], such as those deployed on edge devices, while more complex tasks are offloaded to larger models on the edge server. This architectural separation allows computation to be distributed across edge devices and the edge server in a more natural and efficient manner. As a result, the computational capacity of the edge server is reserved for challenging verification tasks, rather than being consumed by processing simpler tokens from edge devices, unlike in a centralized LLM serving setup.
IV-B Speculative Length vs. throughput and capacity
@@
In speculative decoding, the length of the draft sequence used for verification on the target model is defined as the speculative length, and it affects both token generation throughput and the capacity of SLED, that is, the number of edge devices supported by SLED simultaneously. In this experiment, we manually adjust the speculative length for drafting using LLaMA 1B and 3B models on a Raspberry Pi 5 device, and measure both the device throughput and overall system capacity. As shown in Fig. 5, increasing the speculative length results in lower device throughput but higher system capacity. This inverse relationship between per-device and system-level metrics highlights the importance of selecting an appropriate speculative length to balance the performance of individual edge devices and the system as a whole. On one hand, a longer speculative length reduces token generation on each device, since the drafting throughput remains stable, and a longer speculative length leads to a longer verification period, thereby reducing the response update rate. On the other hand, a longer verification period for individual devices reduces the verification workload on the edge server, allowing it to support more devices concurrently.
IV-C Cost Efficiency and System Throughput
The cost efficiency of token generation in edge computing scenarios is a critical factor and has been considered in various edge inference system designs [29, 30]. In this paper, we compare the proposed SLED system, a centralized serving system, and an all-edge decoding system in terms of both cost efficiency and whole-system token generation rate (WSTGR). To systematically analyze the trade-off between cost and performance, we construct Pareto front visualizations, which highlight the non-dominated configurations that achieve the best balance between monetary cost and system throughput.
In our experiments, the cost and throughput metrics for edge inference scenarios were rigorously computed based on a comprehensive capital expenditure (CAPEX) and operational expenditure (OPEX) model[31]. Specifically, we adopt the widely recognized CPU-hour cost model described by Walker[31] and the edge-compute modeling approach proposed by Eriksson[32]. The CAPEX component was determined by amortizing the purchase price of each edge device (Raspberry Pi 5 and Pi 4B, priced at \$80 [33] and \$75 [34] respectively) over a three-year lifetime, assuming an average device utilization rate of 70%. The OPEX component included electrical consumption calculated from experimentally measured average power draw (8 W for Pi 5 and 5 W for Pi 4B) [35] and industrial electricity rates (0.083 \$/kWh) [36]. Combining these costs, we obtained a unified hourly expense for each device, subsequently normalized by the experimentally measured token generation rates (tokens per second), as shown in Eq. 2. The resulting metric, expressed clearly as dollars per one thousand generated tokens (\$/1K tokens), enabled direct and transparent comparison across different experimental configurations and devices.
$\text{Cost}\,=\frac{1000}{3600\,R}\,\Bigl{(}\frac{P_{\text{device}}}{3\times 8760\times 0.70}\;+\;\frac{P_{\text{avg}}}{1000}\times 0.083\Bigr{)}$ (2)
Figure 6 compares following three deployment strategies along a common cost–performance plane. Specifically, the strategies are: 1) all-Server executes every token‐generation step on a bank of four NVIDIA A100–80 GB GPUs. 2) All-Edge places the same LLaMA draft model on each Raspberry Pi 5, with no server involvement. 3) SLED lets the Raspberry Pi 5 generates draft tokens, which are batch-verified on the A100 cluster of same configuration of centralized scenario. For every strategy we sweep two orthogonal factors: quantization precision (16-, 8-, and 4-bit) and edge-device count $N\in\{1,2,4,8,16\}$. Cost is monetised as dollars per one-thousand verified tokens, while throughput is already-defined WSTGR.
We observe that SLED’s skyline consistently dominates the Pareto frontier, achieving lower cost per 1K verified tokens while sustaining higher overall throughput. For instance, at a budget of approximately \$0.13 per 1K tokens, SLED achieves a throughput of 137 tokens/s—representing a 3.5× improvement over the centralized baseline—while simultaneously reducing cost to just 29% of that incurred by the server-only configuration under comparable conditions. This advantage becomes more pronounced as the number of edge devices increases: whereas the All-Edge throughput saturates due to power constraints (e.g., increasing $N$ from 8 to 16 yields only marginal gains), SLED leverages server-side batching to amortize GPU usage, enabling near-linear throughput scaling beyond $N\geq 4$. Furthermore, quantization universally improves cost efficiency across all schemes by simultaneously reducing energy demand and increasing per-device generation rate. Notably, the 4-bit SLED configuration with 16 devices achieves 138 tokens/s at \$0.13 / 1K tokens, representing a 65% improvement in throughput over the best-performing All-Edge setup, with acceptable additional cost. These results substantiate the claim that SLED enables a superior cost–throughput trade-off, combining local cost efficiency with global throughput.
Overall, the experimental evaluations underscore the significant advantages of SLED in distributed inference scenarios, including throughput, capacity, cost efficiency, showcasing insightful findings in the SLED system to motivate more explorations in the future work.